how do u determine the value of r so the line passing through each pair of points has the given slope. (6,3) , (r,2), m= 1\2
Found 2 solutions by MathPro, ankor@dixie-net.com:
Answer by MathPro(14)

Good Evening,
Let’s start with the equation of a line:
y=mx+c
where m is the slope and c is the intercept on the y axis.
You are given two sets of coordinates and a value for m so let’s insert these values for x, y and m into the equation of a line :
(1) for x=6,y=3 we get 6=(1/2)3+c
(2) for x=r,y=2 we get 2=(1/2)r +c
Now we have two equations and two unknowns. From equation (1) we get c=0 so lets put this into equation (2) to solve for r:
2=(1/2)r + 0
Multiply both sides by 2 to get r =4.
Good Luck!
Selma from Rent-A-Whiz, an online tutoring and homework help website. Just visit www.rentawhiz.com to learn more. Currently, we are offering to solve one math problem for free. Just visit our website, select the homework tab, then click on the math problem submission link and quote coupon code 003 when you submit your problem.

Answer by ankor@dixie-net.com(6693)

how do u determine the value of r so the line passing through each pair of points has the given slope.
(6,3) , (r,2), m= 1\2
:
Use the slope formula: = m
:
=
:
=
Cross multiply
1(r-6) = -1(2)
r - 6 = -2
r = -2 + 6
r = 4
:
Find the equation using the point/slope formula
y - 3 = (x - 6)
y - 3 = x - 3
y = x - 3 + 3
y = x
:
Check the points 6,3 and 4,2 in the equation